OpenStudy (anonymous):

Aidan started taking flying lessons two years ago. He pays an annual flight school fee of $895 and a gasoline fee of $30 each time he takes a flying lesson. The function below can be used to calculate the total yearly flight school fee, f (x), in dollars. f (x) = 895 + 30x, where x represents the number of times he took a flying lesson during the year. Last year he paid $2,065 as a total flight school fee. For how many lessons did he pay a gasoline fee (x)?

jimthompson5910 (jim_thompson5910):

plug in f(x) = 2065, then solve for x f(x) = 895 + 30x 2065 = 895 + 30x 2065 - 895 = 30x 1170 = 30x 30x = 1170 x = 1170/30 x = 39 So he took 39 lessons

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

np